### v1

# On the Statistical Complexity of Estimating VENDI Scores from Empirical Data

Abstract
Reference-free evaluation metrics for generative models have recently been studied in the machine learning community. As a reference-free metric, the VENDI score [1] quantifies the diversity of generative models using matrix-based entropy from information theory. The VENDI score is usually computed through the eigendecomposition of an $n\times n$ kernel matrix for $n$ generated samples. However, due to the high computational cost of eigendecomposition for large $n$, the score is often computed on sample sizes limited to a few tens of thousands. In this paper, we explore the statistical convergence of the VENDI score and demonstrate that for kernel functions with an infinite feature map dimension, the evaluated score for a limited sample size may not converge to the matrix-based entropy statistic. We introduce an alternative statistic called the $t$-truncated VENDI statistic. We show that the existing Nyström method and the FKEA approximation method for the VENDI score will both converge to the defined truncated VENDI statistic given a moderate sample size. We perform several numerical experiments to illustrate the concentration of the empirical VENDI score around the truncated VENDI statistic and discuss how this statistic correlates with the visual diversity of image data.
1 Introduction
The widespread use of generative artificial intelligence (AI) has underscored the need for accurate evaluation of generative models based on the quality and diversity of their generated samples. Users of generative AI often have access to multiple models trained using different algorithms and datasets, necessitating efficient evaluation mechanisms to choose the best model from the available options. The applicability of a model evaluation approach depends on factors such as the required generated sample size, computational cost, and the availability of reference data. Several recent studies on evaluating generative models have introduced assessment methods that relax the requirements for data and computational resources.
Specifically, to enable the evaluation of generative models in settings without reference data, the recent literature has focused on reference-free evaluation scores, which remain applicable in the absence of a reference dataset. The VENDI score [1] is one such reference-free metric that quantifies the diversity of generated data using the entropy of a kernel similarity matrix formulated for the generated samples. As analyzed by [1] and [2], the reference-free assessment of the VENDI score can be interpreted as an unsupervised identification of clusters within the generated data, followed by the calculation of the entropy of the detected cluster variable. Due to its flexibility and adaptability to various domains, the VENDI score has been applied to measure the diversity of samples across different modalities, including image, text, and video data.
Figure 1: Statistical convergence of VENDI score for different sample sizes on ImageNet data: (Left plot) finite-dimension cosine similarity kernel (Right plot) infinite dimension Gaussian kernel with bandwidth $\sigma=40$. DINOv2 embedding (dimension 768) is used in computing the score.
While the VENDI score does not require reference samples, its computational cost increases rapidly with the number of generated samples $n$. In practice, calculating the entropy of the eigenvalues of the $n\times n$ kernel matrix involves performing an eigendecomposition, which requires $O(n^{3})$ computations. As a result, the computational load becomes substantial for a large sample size $n$, and the VENDI score is typically evaluated for sample sizes limited to a few tens of thousands. Consequently, the exact VENDI score, as defined in its original formulation, is usually not computed for sample sizes exceeding 20,000. A key question that arises is whether the VENDI score estimated from empirical data has converged to the underlying statistic, i.e., the limit value as the sample size approaches infinity. The statistical convergence has not been thoroughly investigated for models trained on large-scale datasets, e.g. ImageNet [3], which contain many sample categories and may require a large sample size for proper assessment.
In this work, we study the statistical convergence of the VENDI score and aim to analyze the statistic estimated by the VENDI score for large-scale image, text, and video generative models. We discuss the answer to the convergence question for two types of kernel functions: 1) kernel functions with a finite feature dimension, e.g. the cosine similarity and polynomial kernels, 2) kernel functions with an infinite feature map such as Gaussian (RBF) and Laplace kernels. For kernel functions with a bounded feature dimension $d$, we theoretically and numerically show that a sample size $n=O(d)$ is sufficient to guarantee convergence to the VENDI statistic. For example, the left plot in Figure 1 shows that in the case of the cosine similarity kernel, the VENDI score on $n$ randomly selected ImageNet samples converges as the sample size reaches $5000$, where the feature dimension $d$ (using standard DINOv2 embedding) is 768.
On the other hand, our numerical results for kernel functions with an infinite feature map suggest that in practical scenarios with diverse generative models and datasets, a sample size bounded by 20000 is insufficient for convergence to the VENDI statistic. For example, the right plot in Figure 1 shows the evolution of the VENDI score with Gaussian kernel on the ImageNet data, and the VENDI score keeps growing at a considerable rate even with 20,000 samples. As the costs of computing the exact score will be highly expensive beyond 20,000, it will be hard to empirically estimate the required sample size for convergence to the VENDI statistic.
Given the difference between the VENDI score for $n=O(10^{4})$ and the VENDI statistic (when $n\rightarrow\infty$) in the case of infinite kernel feature dimension, an important question is how to interpret the statistic estimated by the VENDI score when the sample size $n$ is bounded. We attempt to address the question by introducing an alternative statistic, which we call the $t$-truncated VENDI statistic. This statistic is calculated using only the top-$t$ eigenvalues of the kernel covariance matrix, excluding the remaining eigenvalues in the calculation of the VENDI statistic. We prove that a sample size $n=O(t)$ is enough to estimate the $t$-truncated VENDI statistic from $n$ generated samples whether the kernel feature dimension is finite or infinite. This result suggests the affordability of estimating the $t$-truncated VENDI statistic for every kernel function.
To connect the defined $t$-truncated VENDI statistic to existing computation methods for the VENDI score, we show that the standard methods proposed for approximating the VENDI score can be viewed as estimations of the $t$-truncated VENDI statistic. The approximation methods include the Nyström method as applied by [1] and the random Fourier feature-based FKEA method [4] for shift-invariant kernels. We show that both the Nyström and FKEA methods with dimension hyperparameter $t$ provide an estimation of the $t$-truncated VENDI statistic. Our results indicate that the truncated VENDI statistic is implicitly estimated by the computationally efficient methods proposed by [1] and [4].
We perform several numerical experiments to validate the connection between the VENDI scores computed using a bounded $n$ samples and our defined $t$-truncated VENDI statistic. Our numerical results on several image, text, and video datasets and generative models indicate that in the case of a finite-dimension kernel map, the VENDI score efficiently converges to the VENDI statistic. On the other hand, in the case of infinite-dimension Gaussian kernel functions, we numerically observe the growth of the score beyond $n=$10,000. Our numerical results further confirm that the scores computed by Nyström method in [1] and the FKEA method [4] provide tight estimations of the truncated VENDI statistic which could provide a statically affordable substitute for the VENDI statistic in practical scenarios with large-scale datasets. The following is a summary of this work’s contributions:
  • Analyzing the statistical convergence of the reference-free VENDI score under finite and infinite kernel feature maps,
  • Providing numerical evidence on the gap between the VENDI score and VENDI statistic for infinite kernel feature maps with bounded sample size $n$,
  • Introducing the truncated VENDI statistic as a diversity-related statistic for generative models
  • Demonstrating convergence of the Nyström and FKEA methods’ proxy VENDI score to the truncated VENDI statistic.
2 Related Works
Diversity evaluation for generative models Diversity evaluation in generative models can be categorized into two primary types: reference-based and reference-free methods. Reference-based approaches rely on a predefined dataset to assess the diversity of generated data. Metrics such as FID [5] and KID [6] measure the distance between the generated data and the reference, while Recall [7, 8] and Coverage[9] evaluate the extent to which the generative model captures existing modes in the reference dataset. In contrast, the reference-free metrics, VENDI [1][10] and RKE [2], assign diversity scores based on the eigenvalues of a kernel similarity matrix of the generated data. [2]’s results interpret the approach as identifying modes and their frequencies within the generated data followed by entropy calculation for the frequency parameters. In this work, we specifically focus on the statistical convergence of the reference-free VENDI and RKE scores.
Efficient computation of matrix-based entropy. Several strategies have been proposed in the literature to reduce the computational complexity of matrix-based entropy calculations, which involve the computation of matrix eigenvalues—a process that scales cubically with the size of the dataset. [11] propose an efficient algorithm for approximating matrix-based Renyi’s entropy of arbitrary order $\alpha$, which achieves a reduction in computational complexity down to $O(n^{2}sm)$ with $s,m\ll n$. Additionally, kernel matrices can be approximated using low-rank techniques such as incomplete Cholesky decomposition [12, 13] or CUR matrix decompositions [14], which provide substantial computational savings. [10] suggest to leverage Nyström method [15] with $m$ components, which results in $O(nm^{2})$ computational complexity. Further reduction in complexity is possible using Random Fourier Features, as suggested by [4], which allows the computation to scale linearly with $O(n)$ as a function of the dataset size. This work focuses on the latter two methods and which statistics they estimate.
Impact of embedding spaces on diversity evaluation. In our image-related experiments, we used the DinoV2 embedding [16], as [17] demonstrate the alignment of this embedding with human evaluations. We note that the kernel function in the VENDI score can be similarly applied to other embeddings, including the standard InceptionV3[18] and CLIP embeddings [19] as suggested by [20]. Also, in our experiments on text data, we utilized the text-embedding-3-large [21] model, and for the video experiments, we employed the I3D embedding [22]. We use the mentioned embeddings in our experiments, while our theoretical results suggest that the convergence behavior would be similar for other embeddings.
3 Preliminaries
Consider a generative model $\mathcal{G}$ that generates samples from a probability distribution $P_{X}$. To conduct a reference-free evaluation of the model, we suppose the evaluator has access to $n$ independently generated samples from $P_{X}$, denoted by $x_{1},\ldots,x_{n}\in\mathcal{X}$. The assessment task is to estimate the diversity of generative model $\mathcal{G}$ by measuring the variety of the observed generated data, $x_{1},\ldots x_{n}$. In the following subsections, we will discuss kernel functions and their application to define the VENDI diversity score.
3.1 Kernel Functions and Matrices
[…]
$k(x,x^{\prime})\,=\,\bigl{\langle}\phi(x),\phi(x^{\prime})\bigr{\rangle}$ (1)
In this work, we study the evaluation using two types of kernel functions: 1) finite-dimension kernels where dimension $d$ is finite, 2) infinite-dimension kernels where there is no feature map satisfying (1) with a finite $d$ value. A standard example of a finite-dimension kernel is the cosine similarity function where $\phi_{\text{cosine}}(x)=x/\|x\|_{2}$. Also, a widely-used infinite-dimension kernel is the Gaussian (RBF) kernel with bandwidth parameter $\sigma>0$ defined as
$k_{\text{\rm Gaussian}(\sigma)}\bigl{(}x,x^{\prime}\bigr{)}\,:=\,\exp\Bigl{(}-\frac{\bigl{\|}x-x^{\prime}\bigr{\|}^{2}_{2}}{2\sigma^{2}}\Bigr{)}$ (2)
Both the mentioned kernel examples belong to normalized kernels which require $k(x,x)=1$ for every $x$, i.e. the feature map $\phi(x)$ has unit Euclidean norm for every $x$. Given a normalized kernel function, the non-negative eigenvalues of the normalized kernel matrix $\frac{1}{n}K$ for $n$ points $x_{1},\ldots x_{n}$ will sum up to $1$, which means that they form a probability model.
3.2 Matrix-based Entropy Functions and VENDI Score
Given a PSD matrix $A\in\mathbb{R}^{d\times d}$ with a unit trace $\mathrm{Tr}(A)=1$, $A$’s eigenvalues form a probability model. The order-$\alpha$ Renyi entropy of matrix $A$ is defined using the order-$\alpha$ entropy of its eigenvalues as
$H_{\alpha}(A)\,:=\,\frac{1}{1-\alpha}\log\Bigl{(}\sum_{i=1}^{d}\lambda^{\alpha}_{i}\Bigr{)}$ (3)
Specifically, in the case of $\alpha=1$, the above definition reduces to the Shannon entropy of the eigenvalues as $H_{1}(A)\,:=\,\sum_{i=1}^{d}\lambda_{i}\log({1}/{\lambda_{i}})$. Applying the above standard definitions to the normalized kernel matrix $\frac{1}{n}K$, [10] define the order-$\alpha$ VENDI score for samples $x_{1},\ldots,x_{n}$ as
$\mathrm{VENDI}_{\alpha}\bigl{(}x_{1},\ldots,x_{n}\bigr{)}\,:=\,\exp\Bigl{(}H_{\alpha}\bigl{(}\frac{1}{n}K\bigr{)}\Bigr{)}$ (4)
We note that in the case of $\alpha=2$, the definition of $\mathrm{VENDI}_{2}$ is identical to the RKE score proposed by [2]. In this particular case, the score can be formulated using the Frobenius norm of the kernel matrix, which we denote by $\|\cdot\|_{F}$,
$\mathrm{RKE}(x_{1},\ldots,x_{n})\,=\,\mathrm{VENDI}_{2}(x_{1},\ldots,x_{n})\,=\,\Bigl{\|}\frac{1}{n}K\Bigr{\|}^{-2}_{F}$
3.3 Statistical Interpretation of VENDI Score
To derive the VENDI statistic that is supposed to be estimated by the VENDI score, we review the following discussion from [23, 1, 2]. First, note that the normalized kernel matrix $\frac{1}{n}K$, whose eigenvalues are used in the definition of VENDI score, can be written as:
$\frac{1}{n}K=\frac{1}{n}\Phi\Phi^{\top}$ (5)
where $\Phi\in\mathbb{R}^{n\times d}$ is an $n\times d$ matrix whose rows are the feature presentations of samples, i.e., $\phi(x_{1}),\ldots,\phi(x_{n})$. Therefore, the normalized kernel matrix $\frac{1}{n}K$ shares the same non-zero eigenvalues with $\frac{1}{n}\Phi^{\top}\Phi$, in which the multiplication order is flipped. Note that $\frac{1}{n}\Phi^{\top}\Phi$ is defined as the empirical kernel covariance matrix $C_{X}$:
$C_{X}:=\frac{1}{n}\sum_{i=1}^{n}\phi(x_{i})\phi(x_{i})^{\top}=\frac{1}{n}\Phi^{\top}\Phi.$
As a result, the empirical covariance matrix $C_{X}$ and kernel matrix $\frac{1}{n}K$ have the same non-zero eigenvalues and therefore share the same matrix-based entropy value for any order $\alpha$: $H_{\alpha}(\frac{1}{n}K)=H_{\alpha}(C_{X})$. Therefore, if we consider the underlying kernel covariance matrix $\widetilde{C}_{X}=\mathbb{E}_{x\sim P_{X}}\bigl{[}\phi(x)\phi(x)^{\top}\bigr{]}$, we can define the VENDI statistic as follows.
Definition 1.
Given data distribution $P_{X}$, we define the order-$\alpha$ VENDI statistic, ${\mathrm{VENDI}}_{\alpha}(P_{X})$, using the matrix-based entropy of the underlying kernel covariance matrix $\widetilde{C}_{X}=\mathbb{E}_{x\sim P_{X}}\bigl{[}\phi(x)\phi(x)^{\top}\bigr{]}$ as
${\mathrm{VENDI}}_{\alpha}(P_{X})\,:=\,\exp\Bigl{(}H_{\alpha}(\widetilde{C}_{X})\Bigr{)}$ (6)
In the next sections, we study the complexity of estimating the above VENDI statistic from a limited number of samples.
4 Statistical Convergence of Vendi Scores in Finite-Dimension Kernels
Given the definitions of the VENDI score and the VENDI statistic, a relevant question is how many samples are required to accurately estimate the VENDI statistic using the VENDI score. To address this question, we first prove the following concentration bound on the vector of ordered eigenvalues $[\lambda_{1},\ldots,\lambda_{n}]$ of the kernel matrix for a normalized kernel function. We defer the proof of the theoretical results to the Appendix.
Theorem 1.
Consider a normalized kernel function $k$ satisfying $k(x,x)=1$ for every $x\in\mathcal{X}$. Let $\widehat{\boldsymbol{\lambda}}_{n}$ be the vector of sorted eigenvalues of the normalized kernel matrix $\frac{1}{n}K$ for $n$ independent samples $x_{1},\ldots,x_{n}\sim P_{X}$. If we define $\widetilde{\boldsymbol{\lambda}}$ as the vector of sorted eigenvalues of underlying covariance matrix $\widetilde{C}_{X}$, then for every $\delta\geq\exp\bigl{(}(2-n)/8\bigr{)}$, the following inequality holds with probability at least $1-\delta$:
$\bigl{\|}\widehat{\boldsymbol{\lambda}}_{n}-\widetilde{\boldsymbol{\lambda}}\bigr{\|}_{2}\,\leq\,\sqrt{\frac{32\log\bigl{(}2/\delta\bigr{)}}{n}}$
Note that in calculating the subtraction $\widehat{\boldsymbol{\lambda}}_{n}-\widetilde{\boldsymbol{\lambda}}$, we add $|d-n|$ zero entries to the lower-dimension vector, if the dimension of vectors $\widehat{\boldsymbol{\lambda}}_{n}$ and $\widetilde{\boldsymbol{\lambda}}$ do not match.
The above theorem implies the following corollary on a dimension-free convergence guarantee for order-$\alpha$ VENDI score with $\alpha\geq 2$.
Corollary 1.
In the setting of Theorem 1, for every $\alpha\geq 2$ and $\delta\geq\exp\bigl{(}(2-n)/8\bigr{)}$, the following bound holds with probability at least $1-\delta$:
$\Bigl{|}\,\mathrm{VENDI}_{\alpha}\bigl{(}x_{1},\ldots,x_{n}\bigr{)}^{\frac{1-\alpha}{\alpha}}-\mathrm{VENDI}_{\alpha}\bigl{(}P_{X}\bigr{)}^{\frac{1-\alpha}{\alpha}}\,\Bigr{|}\leq\sqrt{\frac{32\log\bigl{(}2/\delta\bigr{)}}{n}}$
On the other hand, if the order $\alpha$ of VENDI score is less than $2$, then Theorem 1 provides a statistical convergence guarantee under a bounded dimension kernel map as shown in the following corollary.
Corollary 2.
In the setting of Theorem 1, consider a finite dimension kernel map where $\mathrm{dim}(\phi)=d<\infty$. Then, for every $21\leq\alpha<$ 2 and $\delta\geq\exp\bigl{(}(2-n)/8\bigr{)}$, the following bound holds with probability at least $1-\delta$:
$\Bigl{|}\,\mathrm{VENDI}_{\alpha}\bigl{(}x_{1},\ldots,x_{n}\bigr{)}^{\frac{1-\alpha}{\alpha}}-\mathrm{VENDI}_{\alpha}\bigl{(}P_{X}\bigr{)}^{\frac{1-\alpha}{\alpha}}\,\Bigr{|}\leq\sqrt{\frac{32d^{2-\alpha}\log\bigl{(}2/\delta\bigr{)}}{n}}$
Therefore, assuming a bounded feature map $d<\infty$, the above result indicates the convergence of the VENDI score to the underlying VENDI statistic given $n=O(d)$ samples. We note that the result is consistent with our numerical observations of the convergence of VENDI score using a finite dimension kernel, e.g. the cosine similarity kernel (Figure 1). Next, we discuss how to extend the above result to an infinite-dimension kernel map by defining the truncated VENDI statistic.
5 Truncated Vendi Statistic and its Estimation via Proxy Kernels
Corollaries 1,2 demonstrate that if the VENDI score order $\alpha$ is greater than $2$ or the kernel feature map dimension $d$ is finite, then the VENDI score can converge to the underlying VENDI statistic with $n=O(d)$ samples. However, the theoretical results do not apply to an order $21\leq\alpha<$ 2 when the kernel map dimension is infinite, e.g. the original order-1 VENDI score [1] with a Gaussian kernel. Our numerical observations indicate that a standard sample size below 20000 could be insufficient for the convergence of order-1 VENDI score (Figure 1). To address this gap, here we define the truncated VENDI statistic and then show the existing kernel approximation algorithms for VENDI score concentrate around this modified statistic.
Definition 2.
Consider data distribution $P$ and its underlying kernel covariance matrix $\widetilde{C}_{X}=\mathbb{E}_{x\sim P}\bigl{[}\phi(x)\phi(x)^{\top}\bigr{]}$. Then, for parameter $t\geq 1$, consider the top-$t$ eigenvalues of $\widetilde{C}_{X}$: $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{t}$. Define $S_{t}=\sum_{i=1}^{t}\lambda_{i}$, and consider the probability sequence $\widetilde{\lambda}_{i}=\lambda_{i}+\frac{1-S_{t}}{t}$ for $i=1,\ldots,t$. Then, we define the order-$\alpha$ $t$-truncated VENDI statistic as
$\mathrm{VENDI}_{\alpha}^{(t)}(P_{X}):=\exp\Bigl{(}\frac{1}{1-\alpha}\log\Bigl{(}\sum_{i=1}^{t}{\widetilde{\lambda}}^{\alpha}_{i}\Bigr{)}\Bigr{)}.$
According to the above definition, we find the probability model with the minimum $\ell_{2}$-norm difference from the $t$-dimensional vector $[\lambda_{1},\ldots,\lambda_{t}]$ including only the top-$t$ eigenvalues. Then, we use the order-$\alpha$ entropy of the probability model to define the order-$\alpha$ $t$-truncated VENDI statistic. Our next result shows that this statistic can be generally estimated using $n=O(t)$ samples.
Theorem 2.
Consider the setting in Theorem 1. Then, for every $\delta\geq\exp((2-n)/8)$, the difference between the $t$-truncated VENDI statistic and its empirical estimation from samples $x_{1},\ldots,x_{n}$ is bounded as follows with probability at least $1-\delta$:
$\Bigl{|}\,\mathrm{VENDI}^{(t)}_{\alpha}\bigl{(}x_{1},\ldots,x_{n}\bigr{)}^{\frac{1-\alpha}{\alpha}}-\mathrm{VENDI}^{(t)}_{\alpha}\bigl{(}P_{X}\bigr{)}^{\frac{1-\alpha}{\alpha}}\,\Bigr{|}\leq\sqrt{\frac{32\max\{1,t^{2-\alpha}\}\log\bigl{(}2/\delta\bigr{)}}{n}}$
As implied by Theorem 2, the $t$-truncated VENDI statistic can be estimated using $O(t)$ samples, i.e. the truncation parameter $t$ plays the role of the bounded dimension of a finite-dimension kernel map. Our next theorem shows that the Nyström method used by [10] and the FKEA method proposed by [4] to reduce the computational complexity of the VENDI score are expected to have a bounded distance to the truncated VENDI statistic.
Theorem 3.
Consider the setting of Theorem 1. (a) Assume that the kernel function is shift-invariant and the FKEA method with $t$ random Fourier features is used to approximate the VENDI score. Then, for every $\delta\geq\exp((2-n)/8)$, with probability at least $1-\delta$:
$\Bigl{|}\mathrm{FKEA}\text{-}\mathrm{VENDI}^{(t)}_{\alpha}\bigl{(}x_{1},\ldots,x_{n}\bigr{)}^{\frac{1-\alpha}{\alpha}}-\mathrm{VENDI}^{(t)}_{\alpha}\bigl{(}P_{X}\bigr{)}^{\frac{1-\alpha}{\alpha}}\Bigr{|}\,\leq\sqrt{\frac{128\max\{1,t^{2-\alpha}\}\log\bigl{(}3/\delta\bigr{)}}{\min\{n,t\}}}$
(b) Assume that the Nyström method is applied with parameter $t$ for approximating the kernel function. Then, if for some $r\geq 1$, the kernel matrix $K$’s $r$th-largest eigenvalue satisfies ${\lambda}_{r}\leq\tau$ and $t\geq r\tau\log(n)$, the following holds with probability at least $1-\delta-2n^{-3}$:
$\Bigl{|}\mathrm{Nystrom}\text{-}\mathrm{VENDI}^{(t)}_{\alpha}\bigl{(}x_{1},\ldots,x_{n}\bigr{)}^{\frac{1-\alpha}{\alpha}}-\mathrm{VENDI}^{(t)}_{\alpha}\bigl{(}P_{X}\bigr{)}^{\frac{1-\alpha}{\alpha}}\Bigr{|}\leq\mathcal{O}\Bigl{(}\sqrt{\frac{\max\{1,t^{2-\alpha}\}\log\bigl{(}2/\delta\bigr{)}t\tau^{2}\log(n)^{2}}{n}}\Bigr{)}$
Theorem 3 shows that the application of the fast kernel approximation algorithms in FKEA and Nyström with parameter $t$ will provide an estimation of the $t$-truncated VENDI statistic, connecting the defined statistic to the existing algorithms for computing the VENDI score.
6 Numerical Results
We evaluated the convergence of the VENDI score, the truncated VENDI score, and the proxy VENDI scores using the Nyström method and FKEA in our numerical experiments. We provide a comparative analysis of these scores across different data types and models, including image, text, and video. In our experiments, we considered the cosine similarity kernel as a standard kernel function with a finite-dimension map and the Gaussian (RBF) kernel as a kernel function with an infinite-dimension feature map. In the experiments with Gaussian kernels, we matched the kernel bandwidth parameter with those chosen by [2, 4] for the same datasets. We used 20,000 number of samples per score computation, consistent with standard practice in the literature. To investigate how computation-cutting methods compare to each other, in the experiments we matched the truncation parameter $t$ of our defined $t$-truncated VENDI score with the Nyström method’s hyperparameter on the number of randomly selected rows of kernel matrix and the FKEA’s hyperparameter of the number of random Fourier features. The VENDI and FKEA implementations were adopted from the corresponding references’ GitHub webpages, while the Nyström method was adopted from the scikit-learn Python package.
6.1 Convergence Analysis of VENDI scores
To assess the convergence of the discussed VENDI scores, we conducted experiments on four datasets including ImageNet and FFHQ[24] image datasets, a synthetic text dataset with 400k paragraphs generated by GPT-4 about 100 randomly selected countries, and the Kinetics video dataset [25]. Our results, presented in Figure LABEL:VENDI_convergence, show that for the finite-dimension cosine similarity kernel the VENDI score converges rapidly to the underlying value and the proxy versions including truncated and Nyström VENDI scores were almost identical to the original VENDI score. This observation is consistent with our theoretical results on the convergence of VENDI scores under finite-dimension kernel maps. On the other hand, in the case of infinite dimension Gaussian kernel, we observed that the $\mathrm{VENDI}_{1}$ score did not converge using 20k samples and the score value kept growing with a considerable rate. However, the $t$-truncated VENDI score with $t=10000$ converged to its underlying statistic shortly after 10000 samples were used. Consistent with our theoretical result, the proxy Nyström and FKEA estimated scores with their rank hyperparameter matched with $t$ also converged to the limit of the truncated VENDI scores. The numerical results show the connection between the truncated VENDI statistic and the existing kernel-based algorithms for approximating the VENDI score.
6.2 Correlation between the truncated VENDI statistic and diversity of data
[…]
7 Conclusion
In this work, we investigated the statistical convergence properties of VENDI diversity scores estimated from empirical samples. We highlighted that, due to the high computational complexity of the score for datasets larger than a few tens of thousands of generated data points, the score is often calculated using sample sizes below 20,000. We demonstrated that such limited sample sizes do not pose a problem for statistical convergence as long as the kernel feature dimension is bounded. Conversely, our numerical results showed a lack of convergence to the VENDI statistic when using an infinite-dimensional kernel map, such as the Gaussian kernel. To address this issue, we introduced the truncated VENDI statistic as an alternative target statistic. We further showed that existing methods, such as Nyström and FKEA, for approximating VENDI scores concentrate well around this truncated VENDI statistic. A promising direction for future research is to explore the relationship between other kernel approximation techniques and the truncated VENDI statistic. Moreover, a comprehensive analysis of the computational-statistical trade-offs involved in estimating the VENDI score remains an open question for future work.
Appendix A Proofs
A.1 Proof of Theorem 1
To prove the theorem, we will use the following lemma followed from [28, 29].
Lemma 1 (Vector Bernstein Inequality [28, 29]).
Suppose that $z_{1},\ldots,z_{n}$ are independent and identically distributed random vectors with zero mean $\mathbb{E}[z_{i}]=\mathbf{0}$ and bounded $\ell_{2}$-norm $\|z_{i}\|_{2}\leq c$. Then, for every $0\leq\epsilon\leq c$, the following holds
$\mathbb{P}\biggl{(}\,\Bigl{\|}\frac{1}{n}\sum_{i=1}^{n}z_{i}\Bigr{\|}_{2}\geq\epsilon\,\biggr{)}\,\leq\,\exp\Bigl{(}-\frac{n\epsilon^{2}}{8c^{2}}+\frac{1}{4}\Bigr{)}$
We apply the above Vector Bernstein Inequality to the random vectors $\phi(x_{1})\otimes\phi(x_{1}),\ldots,\phi(x_{1})\otimes\phi(x_{1})$ where $\otimes$ denotes the Kronecker product. To do this, we define vector $v_{i}=\phi(x_{i})\otimes\phi(x_{i})-\mathbb{E}_{x\sim P}\bigl{[}\phi(x)\otimes\phi(x)\bigr{]}$ for every $i$. Note that $v_{i}$ is, by definition, a zero-mean vector and also for every $x$ we have the following for the normalized kernel function $k$:
$\bigl{\|}\phi(x)\otimes\phi(x)\bigr{\|}^{2}_{2}=\bigl{\|}\phi(x)\bigr{\|}^{2}_{2}\cdot\bigl{\|}\phi(x)\bigr{\|}^{2}_{2}=k(x,x)\cdot k(x,x)=1$
Then, the triangle inequality implies that
$\bigl{\|}v_{i}\bigr{\|}_{2}\leq\bigl{\|}\phi(x_{i})\otimes\phi(x_{i})\bigr{\|}_{2}+\bigl{\|}\mathbb{E}_{x\sim P}\bigl{[}\phi(x)\otimes\phi(x)\bigr{]}\bigr{\|}_{2}\leq\bigl{\|}\phi(x_{i})\otimes\phi(x_{i})\bigr{\|}_{2}+\mathbb{E}_{x\sim P}\bigl{[}\bigl{\|}\phi(x)\otimes\phi(x)\bigr{\|}_{2}\bigr{]}=2$
As a result, the Vector Bernstein Inequality leads to the following for every $0\leq\epsilon\leq 2$:
$\mathbb{P}\Bigl{(}\Bigl{\|}\frac{1}{n}\sum_{i=1}^{n}\phi(x_{i})\otimes\phi(x_{i})-\mathbb{E}_{x\sim P}\bigl{[}\phi(x)\otimes\phi(x)\bigr{]}\Bigr{\|}_{2}\geq\epsilon\Bigr{)}\,\leq\,\exp\Bigl{(}\frac{8-n\epsilon^{2}}{32}\Bigr{)}$
On the other hand, note that $\phi(x)\otimes\phi(x)$ is the vectorized version of rank-1 matrix $\phi(x)\phi(x)^{\top}$, which shows that the above inequality is equivalent to
$\mathbb{P}\Bigl{(}\Bigl{\|}\frac{1}{n}\sum_{i=1}^{n}\bigl{[}\phi(x_{i})\phi(x_{i})^{\top}\bigr{]}-\mathbb{E}_{x\sim P}\bigl{[}\phi(x)\phi(x)^{\top}\bigr{]}\Bigr{\|}_{F}\geq\epsilon\Bigr{)}\,\leq\,\exp\Bigl{(}\frac{8-n\epsilon^{2}}{32}\Bigr{)}$
[…]
Subsequently, we can apply the Hoffman-Wielandt inequality which shows that for the sorted eigenvalue vectors of $C_{X}$ (denoted by $\widehat{\boldsymbol{\lambda}}_{n}$ in the theorem) and $\widetilde{C}_{X}$ (denoted by $\widetilde{\boldsymbol{\lambda}}$ in the theorem) we will have $\|\widehat{\boldsymbol{\lambda}}_{n}-\widetilde{\boldsymbol{\lambda}}\|_{2}\leq\|C_{X}-\widetilde{C}_{X}\|_{F}$, which together with the previous inequality leads to
$\mathbb{P}\Bigl{(}\bigl{\|}\widehat{\boldsymbol{\lambda}}_{n}-\widetilde{\boldsymbol{\lambda}}\bigr{\|}_{2}\geq\epsilon\Bigr{)}\,\leq\,\exp\Bigl{(}\frac{8-n\epsilon^{2}}{32}\Bigr{)}$
If we define $\delta=\exp\bigl{(}(8-n\epsilon^{2})/32\bigr{)}$ that implies $\epsilon\leq\sqrt{\frac{32\log(2/\delta)}{n}}$, we obtain the following for every $\delta\geq\exp\bigl{(}(2-n)/8\bigr{)}$ (since we suppose $0\leq\epsilon\leq 2$)
$\mathbb{P}\Bigl{(}\bigl{\|}\widehat{\boldsymbol{\lambda}}_{n}-\widetilde{\boldsymbol{\lambda}}\bigr{\|}_{2}\geq\sqrt{\frac{32\log(2/\delta)}{n}}\Bigr{)}\,\leq\,\delta$
[…]
which completes the proof.
A.2 Proof of Corollary 1
Considering the $\alpha$-norm definition $\|\mathbf{v}\|_{\alpha}=\bigl{(}\sum_{i=1}^{d}|v_{i}|^{\alpha}\bigr{)}^{1/\alpha}$, we can rewrite the order-$\alpha$ VENDI definition as
$\mathrm{VENDI}_{\alpha}(x_{1},\ldots,x_{n})=\bigl{\|}\widehat{\boldsymbol{\lambda}}_{n}\bigr{\|}^{\frac{\alpha}{1-\alpha}}_{\alpha}\quad\Longleftrightarrow\quad\mathrm{VENDI}_{\alpha}(x_{1},\ldots,x_{n})^{\frac{1-\alpha}{\alpha}}=\bigl{\|}\widehat{\boldsymbol{\lambda}}_{n}\bigr{\|}_{\alpha}$
where $\widehat{\boldsymbol{\lambda}}_{n}$ is defined in Theorem 1. Similarly, given the definition of $\widetilde{\boldsymbol{\lambda}}$ we can write
$\mathrm{VENDI}_{\alpha}(P_{x})^{\frac{1-\alpha}{\alpha}}=\bigl{\|}\widetilde{\boldsymbol{\lambda}}\bigr{\|}_{\alpha}$
Therefore, for every $\alpha\geq 2$, the following hold due to the triangle inequality:
$\Bigl{|}\mathrm{VENDI}_{\alpha}(x_{1},\ldots,x_{n})^{\frac{1-\alpha}{\alpha}}-\mathrm{VENDI}_{\alpha}(P_{x})^{\frac{1-\alpha}{\alpha}}\Bigr{|}\,$ $=\,\Bigl{|}\bigl{\|}\widehat{\boldsymbol{\lambda}}_{n}\bigr{\|}_{\alpha}-\bigl{\|}\widetilde{\boldsymbol{\lambda}}\bigr{\|}_{\alpha}\Bigr{|}$
$\leq\,\bigl{\|}\widehat{\boldsymbol{\lambda}}_{n}-\widetilde{\boldsymbol{\lambda}}\bigr{\|}_{\alpha}$
$\leq\,\bigl{\|}\widehat{\boldsymbol{\lambda}}_{n}-\widetilde{\boldsymbol{\lambda}}\bigr{\|}_{2}.$
As a result, Theorem 1 shows that for every $\alpha\geq 2$ and $\delta\geq\exp((2-n)/8)$, the following holds with probability at least $1-\delta$
$\Bigl{|}\mathrm{VENDI}_{\alpha}(x_{1},\ldots,x_{n})^{\frac{1-\alpha}{\alpha}}-\mathrm{VENDI}_{\alpha}(P_{x})^{\frac{1-\alpha}{\alpha}}\Bigr{|}\,\leq\,\sqrt{\frac{32\log(2/\delta)}{n}}$
A.3 Proof of Corollary 2
Given $21\leq\alpha<$ 2, the inequality $\|v\|_{\alpha}\leq d^{\frac{2-\alpha}{2}}\|v\|_{2}$ holds for every $d$-dimensional vector $v\in\mathbb{R}^{d}$. Therefore, we can repeat the proof of Corollary 1 to show the following for every $21\leq\alpha<$ 2
$\Bigl{|}\mathrm{VENDI}_{\alpha}(x_{1},\ldots,x_{n})^{\frac{1-\alpha}{\alpha}}-\mathrm{VENDI}_{\alpha}(P_{x})^{\frac{1-\alpha}{\alpha}}\Bigr{|}\,$ $=\,\Bigl{|}\bigl{\|}\widehat{\boldsymbol{\lambda}}_{n}\bigr{\|}_{\alpha}-\bigl{\|}\widetilde{\boldsymbol{\lambda}}\bigr{\|}_{\alpha}\Bigr{|}$
$\leq\,\bigl{\|}\widehat{\boldsymbol{\lambda}}_{n}-\widetilde{\boldsymbol{\lambda}}\bigr{\|}_{\alpha}$
$\leq\,d^{\frac{2-\alpha}{2}}\bigl{\|}\widehat{\boldsymbol{\lambda}}_{n}-\widetilde{\boldsymbol{\lambda}}\bigr{\|}_{2}.$
Consequently, Theorem 1 impies that for every $21\leq\alpha<$ 2 and $\delta\geq\exp((2-n)/8)$, the following holds with probability at least $1-\delta$
$\Bigl{|}\mathrm{VENDI}_{\alpha}(x_{1},\ldots,x_{n})^{\frac{1-\alpha}{\alpha}}-\mathrm{VENDI}_{\alpha}(P_{x})^{\frac{1-\alpha}{\alpha}}\Bigr{|}\,\leq\,d^{\frac{2-\alpha}{2}}\sqrt{\frac{32\log(2/\delta)}{n}}\,=\,\sqrt{\frac{32d^{2-\alpha}\log(2/\delta)}{n}}$
A.4 Proof of Theorem 2
We begin by proving the following lemma showing that the eigenvalues used in the definition of the $t$-truncated VENDI score are the projection of the original eigenvalues onto a $t$-dimensional probability simplex.
Lemma 2.
Consider $\mathbf{v}\in[0,1]^{d}$ that satisfies $\mathbf{1}^{\top}\mathbf{v}=1$. i.e., the sum of $\mathbf{v}$’s entries equals $1$. Given integer $1\leq t\leq d$, define vector $\mathbf{v}^{(t)}\in[0,1]^{d}$ whose last $d-t$ entries are 00, i.e., $v^{(t)}_{i}=0$ for $t+1\leq i\leq d$, and its first $t$ entries are defined as $v^{(t)}_{j}=v_{j}+\frac{1-S_{t}}{t}$ where $S_{t}=v_{1}+\cdots+v_{t}$. Then, $\mathbf{v}^{(t)}$ is the projection of $\mathbf{v}$ onto the following simplex set and has the minimum $\ell_{2}$-norm distance to this set
$\Delta_{t}:=\Bigl{\{}\mathbf{u}\in[0,1]^{d}:\;v_{i}=0\>\text{\rm for all}\>t+1\leq i\leq d,\;\;\;\sum_{i=1}^{t}v_{i}=1\Bigr{\}}.$
Proof.
To prove the lemma, first note that $\mathbf{v}^{(t)}\in\Delta_{t}$, i.e. its first $t$ entries are non-negative and add up to $1$, and also its last $d-t$ entries are zero. Then, consider the projection problem discussed in the lemma:
$\min_{\mathbf{u}\in\mathbb{R}^{t}}\qquad\;\;\>\sum_{i=1}^{t}\bigl{(}u_{i}-v_{i}\bigr{)}^{2}$
$\text{\rm subject to}\quad u_{i}\geq 0,\;\;\text{\rm for all}\>i$
$\qquad\qquad\quad\sum_{i=1}^{t}u_{i}=1$
Then, since we know from the assumptions that $v_{i}\geq 0$ and $\sum_{i=1}^{t}v_{i}\leq 1$, the discussed $\mathbf{u}^{*}\in\mathbb{R}^{t}$ where $u^{*}_{i}=v_{i}+(1-S_{t})/t$ together with Lagrangian coefficients $\mu_{i}=0$ (for inequality constraint $u_{i}\geq 0$) and $\lambda=(1-S_{t})/t$ (for equality constraint) satisfy the KKT conditions. The primal and dual feasibility conditions as well as the complementary slackness clearly hold for these selection of primal and dual variables. Also, the KKT stationarity condition is satisfied as for every $i$ we have $u^{*}_{i}-v_{i}-\lambda-\mu_{i}=0$. Since the optimization problem is a convex optimization task with affine constraints, the KKT conditions are sufficient for optimaility which proves the lemma. ∎
Based on the above lemma, the eigenvalues $\widehat{\boldsymbol{\lambda}}_{n}^{(t)}$ used to calculate the $t$-truncated VENDI score $\mathrm{VENDI}^{(t)}_{\alpha}(x_{1},\ldots,x_{n})$ are the projections of the top-$t$ eigenvalues in $\widehat{\boldsymbol{\lambda}}_{n}$ for the original score $\mathrm{VENDI}_{\alpha}(x_{1},\ldots,x_{n})$ onto the $t$-simplex subset of $\mathbb{R}^{d}$ according to the $\ell_{2}$-norm. Similarly, the eigenvalues $\widetilde{\boldsymbol{\lambda}}_{n}^{(t)}$ used to calculate the $t$-truncated VENDI statistic $\mathrm{VENDI}^{(t)}_{\alpha}(P_{X})$ are the projections of the top-$t$ eigenvalues in $\widetilde{\boldsymbol{\lambda}}$ for the original VENDI statistic $\mathrm{VENDI}_{\alpha}(P_{x})$ onto the $t$-simplex subset of $\mathbb{R}^{d}$.
Since $\ell_{2}$-norm is a Hilbert space norm and the $t$-simplex subset $\Delta_{t}$ is a convex set, we know from the convex analysis that the $\ell_{2}$-distance between the projected points $\widehat{\boldsymbol{\lambda}}_{n}^{(t)}$ and $\widetilde{\boldsymbol{\lambda}}^{(t)}$ is upper-bounded by the $\ell_{2}$-distance between the original points $\widehat{\boldsymbol{\lambda}}_{n}$ and $\widetilde{\boldsymbol{\lambda}}$. As a result, Theorem 1 implies that
[…]
However, note that the eigenvalue vectors $\widehat{\boldsymbol{\lambda}}^{(t)}_{n}$ and $\widetilde{\boldsymbol{\lambda}}^{(t)}$ can be analyzed in a bounded $t$-dimensional space as their entries after index $t+1$ are zero. Therefore, we can apply the proof of Corollary 2 to show that for every $21\leq\alpha<$ 2 and $\delta\geq\exp((2-n)/8)$, the following holds with probability at least $1-\delta$
$\Bigl{|}\mathrm{VENDI}_{\alpha}(x_{1},\ldots,x_{n})^{\frac{1-\alpha}{\alpha}}-\mathrm{VENDI}_{\alpha}(P_{x})^{\frac{1-\alpha}{\alpha}}\Bigr{|}\,\leq\,\sqrt{\frac{32t^{2-\alpha}\log(2/\delta)}{n}}$
To extend the result to a general $\alpha>1$, we reach the following inequality covering the above result as well as the result of Corollary 1 in one inequality
$\Bigl{|}\mathrm{VENDI}_{\alpha}(x_{1},\ldots,x_{n})^{\frac{1-\alpha}{\alpha}}-\mathrm{VENDI}_{\alpha}(P_{x})^{\frac{1-\alpha}{\alpha}}\Bigr{|}\,\leq\,\sqrt{\frac{32\max\{1,t^{2-\alpha}\}\log(2/\delta)}{n}}$
A.5 Proof of Theorem 3
Proof of Part (a). As defined by [4], the FKEA method uses the eigenvalues of $t$ random Fourier frequencies $\omega_{1},\ldots,\omega_{t}$ where for each $\omega_{i}$ they consider two features $\cos(\omega_{i}^{\top}x)$ and $\sin(\omega_{i}^{\top}x)$. Following the definitions, it can be seen that $k(x,x^{\prime})=\mathbb{E}_{\omega\sim p_{\omega}}\bigl{[}\cos(\omega^{\top}(x-x^{\prime}))\bigr{]}$ which is approximated by FKEA as $\frac{1}{t}\sum_{i=1}^{t}\cos(\omega_{i}^{\top}(x-x^{\prime}))$. Therefore, if we define kernel matrix $K_{i}$ as the kernel matrix for $k_{i}(x,x^{\prime})=\cos(\omega_{i}^{\top}(x-x^{\prime}))$, then we will have
$\frac{1}{n}K^{\mathrm{FKEA}(t)}=\frac{1}{t}\sum_{i=1}^{t}\frac{1}{n}K_{i}$
where $\mathbb{E}_{\omega_{i}\sim p_{\omega}}[\frac{1}{n}K_{i}]=\frac{1}{n}K$.
On the other hand, we note that $\|\frac{1}{n}K\|_{F}\leq 1$ holds as the kernel function is normalized and hence $|k(x,x^{\prime})|\leq 1$. Since the Frobenius norm is the $\ell_{2}$-norm of the vectorized version of the matrix, we can apply Vector Bernstein inequality in Lemma 1 to show that for every $0\leq\epsilon\leq 2$:
$\mathbb{P}\Bigl{(}\Bigl{\|}\frac{1}{t}\sum_{i=1}^{t}\bigl{[}\frac{1}{n}K_{i}\bigr{]}-\frac{1}{n}K\Bigr{\|}_{F}\geq\epsilon\Bigr{)}\,\leq\,\exp\Bigl{(}\frac{8-t\epsilon^{2}}{32}\Bigr{)}$
[…]
Then, we apply the Hoffman-Wielandt inequality to show that for the sorted eigenvalue vectors of $\frac{1}{n}K$ (denoted by $\widehat{\boldsymbol{\lambda}}_{n}$) and $\frac{1}{n}K^{\mathrm{FKEA}(t)}$ (denoted by ${\boldsymbol{\lambda}}^{\mathrm{FKEA}(t)}$) we will have $\|\widehat{\boldsymbol{\lambda}}_{n}-{\boldsymbol{\lambda}}^{\mathrm{FKEA}(t)}\|_{2}\leq\|\frac{1}{n}K^{\mathrm{FKEA}(t)}-\frac{1}{n}K\|_{F}$, which together with the previous inequality leads to
$\mathbb{P}\Bigl{(}\Bigl{\|}\widehat{\boldsymbol{\lambda}}_{n}-{\boldsymbol{\lambda}}^{\mathrm{FKEA}(t)}\Bigr{\|}_{2}\geq\epsilon\Bigr{)}\,\leq\,\exp\Bigl{(}\frac{8-t\epsilon^{2}}{32}\Bigr{)}$
Furthermore, as we shown in the proof of Theorem 1 for every $0\leq\gamma\leq 2$
$\mathbb{P}\Bigl{(}\bigl{\|}\widehat{\boldsymbol{\lambda}}_{n}-\widetilde{\boldsymbol{\lambda}}\bigr{\|}_{2}\geq\gamma\Bigr{)}\,\leq\,\exp\Bigl{(}\frac{8-n\gamma^{2}}{32}\Bigr{)}$
which, by applying the union bound for $\gamma=\epsilon$, together with the previous inequality shows that
$\mathbb{P}\Bigl{(}\Bigl{\|}\widetilde{\boldsymbol{\lambda}}-{\boldsymbol{\lambda}}^{\mathrm{FKEA}(t)}\Bigr{\|}_{2}\geq 2\epsilon\Bigr{)}\,$ $\leq\,\exp\Bigl{(}\frac{8-t\epsilon^{2}}{32}\Bigr{)}+\exp\Bigl{(}\frac{8-n\epsilon^{2}}{32}\Bigr{)}$
 $\leq\,2\exp\Bigl{(}\frac{8-\min\{n,t\}\epsilon^{2}}{32}\Bigr{)}$
Therefore, Lemma 2 implies that
$\mathbb{P}\Bigl{(}\Bigl{\|}\widetilde{\boldsymbol{\lambda}}^{(t)}-{\boldsymbol{\lambda}}^{\mathrm{FKEA}(t)}\Bigr{\|}_{2}\geq\epsilon\Bigr{)}\,$  $\leq\,2\exp\Bigl{(}\frac{32-\min\{n,t\}\epsilon^{2}}{128}\Bigr{)}$
If we define $\delta=2\exp\bigl{(}\frac{32-\min\{n,t\}\epsilon^{2}}{128}\bigr{)}$, implying that $\epsilon\leq\sqrt{\frac{128\log(3/\delta)}{\min\{n,t\}}}$, then the above inequality shows that
$\mathbb{P}\Bigl{(}\Bigl{\|}\widetilde{\boldsymbol{\lambda}}^{(t)}-{\boldsymbol{\lambda}}^{\mathrm{FKEA}(t)}\Bigr{\|}_{2}\leq\sqrt{\frac{128\log(3/\delta)}{\min\{n,t\}}}\Bigr{)}\,$ $\geq 1-\delta$
Therefore, if we follow the same steps of the proof of Theorem 2, we can show
$\Bigl{|}\mathrm{FKEA}\text{-}\mathrm{VENDI}^{(t)}_{\alpha}(x_{1},\ldots,x_{n})^{\frac{1-\alpha}{\alpha}}-\mathrm{VENDI}^{(t)}_{\alpha}(P_{x})^{\frac{1-\alpha}{\alpha}}\Bigr{|}\,\leq\,\sqrt{\frac{128\max\{1,t^{2-\alpha}\}\log(3/\delta)}{\min\{n,t\}}}$
Proof of Part (b). To show this theorem, we use Theorem 3 from [30], which shows that if the $r$th largest eigenvalue of the kernel matrix $\frac{1}{n}K$ satisfies $\lambda_{r}\leq\frac{\tau}{n}$, then given $t\geq Cr\log(n)$ ($C$ is a universal constant), the following spectral norm bound will hold with probability $1-\frac{2}{n^{3}}$:
$\bigl{\|}\frac{1}{n}K-\frac{1}{n}K^{\mathrm{Nystrom(t)}}\bigr{\|}_{sp}\leq\mathcal{O}\Bigl{(}\frac{\tau\log(n)}{\sqrt{nt}}\Bigr{)}.$
Therefore, Weyl’s inequality implies the following for the vector of sorted eigenvalues of $\frac{1}{n}K$, i.e. $\widehat{\boldsymbol{\lambda}}_{n}$, and that of $\frac{1}{n}K^{\mathrm{Nystrom(t)}}$, i.e., ${\boldsymbol{\lambda}}^{\mathrm{Nystrom(t)}}$,
$\bigl{\|}\widehat{\boldsymbol{\lambda}}_{n}-{\boldsymbol{\lambda}}^{\mathrm{Nystrom(t)}}\bigr{\|}_{\infty}\leq\mathcal{O}\Bigl{(}\frac{\tau\log(n)}{\sqrt{nt}}\Bigr{)}.$
As a result, considering the subvectors $\widehat{\boldsymbol{\lambda}}_{n}[1:t]$ and ${\boldsymbol{\lambda}}^{\mathrm{Nystrom(t)}}[1:t]$ with the first $t$ entries of the vectors, we will have:
$\bigl{\|}\widehat{\boldsymbol{\lambda}}_{n}[1:t]-{\boldsymbol{\lambda}}^{\mathrm{Nystrom(t)}}[1:t]\bigr{\|}_{\infty}\leq\mathcal{O}\Bigl{(}\frac{\tau\log(n)}{\sqrt{nt}}\Bigr{)}\quad\Longrightarrow\quad\bigl{\|}\widehat{\boldsymbol{\lambda}}_{n}[1:t]-{\boldsymbol{\lambda}}^{\mathrm{Nystrom(t)}}[1:t]\bigr{\|}_{2}\leq\mathcal{O}\Bigl{(}\tau\log(n)\sqrt{\frac{t}{n}}\Bigr{)}$
Noting that the non-zero entries of ${\boldsymbol{\lambda}}^{\mathrm{Nystrom(t)}}$ are all included in the first-$t$ elements, we can apply Lemma 2 which shows that with probability $1-2n^{-3}$ we have
$\Bigl{\|}\widehat{\boldsymbol{\lambda}}^{(t)}_{n}-{\boldsymbol{\lambda}}^{\mathrm{Nystrom(t)}}\Bigr{\|}_{2}\leq\mathcal{O}\Bigl{(}\tau\log(n)\sqrt{\frac{t}{n}}\Bigr{)}$
Also, in the proof of Theorem 2, we showed that
[…]
Combining the above inequalities using a union bound, shows that with probability at least $1-\delta-2n^{-3}$ we have
$\Bigl{\|}{\boldsymbol{\lambda}}^{\mathrm{Nystrom(t)}}-\widetilde{\boldsymbol{\lambda}}^{(t)}\Bigr{\|}_{2}\>$ $\leq\>\Bigl{\|}{\boldsymbol{\lambda}}^{\mathrm{Nystrom(t)}}-\widehat{\boldsymbol{\lambda}}^{(t)}_{n}\Bigr{\|}_{2}+\Bigl{\|}\widehat{\boldsymbol{\lambda}}^{(t)}_{n}-\widetilde{\boldsymbol{\lambda}}^{(t)}\Bigr{\|}_{2}$
$\leq\>\sqrt{\frac{32\log(2/\delta)}{n}}+\mathcal{O}\Bigl{(}\tau\log(n)\sqrt{\frac{t}{n}}\Bigr{)}$
$=\>\mathcal{O}\Bigl{(}\sqrt{\frac{\log(2/\delta)+t\log(n)^{2}\tau^{2}}{n}}\Bigr{)}$
Hence, repeating the final steps in the proof of Theorem 2, we can prove
$\Bigl{|}\mathrm{Nystrom}\text{-}\mathrm{VENDI}^{(t)}_{\alpha}(x_{1},\ldots,x_{n})^{\frac{1-\alpha}{\alpha}}-\mathrm{VENDI}^{(t)}_{\alpha}(P_{x})^{\frac{1-\alpha}{\alpha}}\Bigr{|}\,\leq\,\mathcal{O}\Bigl{(}\sqrt{\frac{\max\{t^{2-\alpha},1\}\bigl{(}\log(2/\delta)+t\log(n)^{2}\tau^{2}\bigr{)}}{n}}\Bigr{)}$
Appendix B Additional Numerical Results
Figure 4: Diversity evaluation of VENDI scores on StyleGAN-XL generated ImageNet dataset with varying truncation parameter $\psi$. The setting is based on DinoV2 embedding and bandwidth $\sigma=30$
In this section, we present supplementary results concerning the evaluation of diversity and the convergence behavior of different variants of the VENDI score. We extend the convergence experiments discussed in the main text to include the truncated StyleGAN3-t FFHQ dataset (Figure LABEL:VENDI_stylegan3_convergence) and the StyleGAN-XL ImageNet dataset (Figure LABEL:VENDI_styleganxl_convergence). Furthermore, we demonstrate that the truncated VENDI statistic effectively captures the diversity characteristics across various data modalities. Specifically, we conducted similar experiments as shown in Figures LABEL:VENDI_imagenet_truncation and LABEL:VENDI_ffhq_truncation on text data (Figure LABEL:VENDI_countries_diversity) and video data (Figure LABEL:VENDI_video_diversity), showcasing the applicability of the metric across different domains.
Figure 6: The diagram outlining an intuition behind a kernel bandwidth $\sigma$ selection in diversity evaluation.
We observe in Figure LABEL:VENDI_stylegan3_convergence that the convergence behavior is illustrated across various values of $\psi$. The results indicate that, for a fixed bandwidth $\sigma$, the truncated, Nyström, and FKEA variants of the VENDI score converge to the truncated VENDI statistic. As demonstrated in Figure LABEL:VENDI_ffhq_truncation of the main text, this truncated VENDI statistic effectively captures the diversity characteristics inherent in the underlying dataset.
We note that in presence of incremental changes to the diversity of the dataset, finite-dimensional kernels, such as cosine similarity kernel, remain relatively constant. This effect is illustrated in Figure LABEL:VENDI_styleganxl_convergence, where increase in truncation factor $\psi$ results in incremental change in diversity. This is one of the cases where infinite-dimensional kernel maps with a sensitivity (bandwidth) parameter $\sigma$ are useful in controlling how responsive the method should be to the change in diversity.
B.1 Approximation error of VENDI scores
In this section, we demonstrate the numerical stability of all VENDI score variants. Figure LABEL:error_bounds presents the standard deviation (expressed as a percentage) across varying sample sizes. The results indicate that, for each sample size $n$, the metric exhibits relatively low variance when computed across multiple sample sets drawn from the same underlying distribution, highlighting the robustness of the VENDI score computation. We present the results for ImageNet and FFHQ datasets.
B.2 Bandwidth $\sigma$ Selection
In our experiments, we select the Gaussian kernel bandwidth, $\sigma$, to ensure that the Vendi metric effectively distinguishes the inherent modes within the dataset. The kernel bandwidth directly controls the sensitivity of the metric to the underlying data clusters. As illustrated in Figure 6, varying $\sigma$ significantly impacts the diversity computation on the ImageNet dataset. A smaller bandwidth (e.g., $\sigma=20,30$) results in the metric treating redundant samples as distinct modes, artificially inflating the number of clusters, which in turn slows down the convergence of the metric. On the other hand, large bandwidth results in instant convergence of the metric, i.e. in $\sigma=60$ $n=100$ and $n=1000$ have almost the same amount of diversity.